# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ VI.3 Quantization
which is the quantum version of the equation of motion that results from the Lagrangian (127). A similar equation can be derived for a conventional Floquet system, where the frequency is fixed by the external driving provided by the periodic Hamiltonian $\hat{H}(\phi)$. Once the dynamics in the $(t,\phi)$ formalism is derived, the actual field operator is retrieved as $\hat{\Psi}(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\hat{\Psi}(\mathbf{x},\phi_{0}+\omega t,t)$. We stress that this equation of motion is exact because it is just a reparametrization of the standard Heisenberg equation of motion. Consequently, the $(t,\phi)$ formalism can be applied along with the Floquet thermodynamics to an arbitrary many-body system.
In the specific case of a condensate, after expanding the field operator around the spontaneous Floquet state, $\hat{\Psi}(\mathbf{x},\phi,t)=\Psi_{0}(\mathbf{x},\phi)+\hat{\varphi}(\mathbf{x},\phi,t)$, we obtain at linear order from Eq. (140) that
$i\partial_{t}\hat{\Phi}=[M_{0}(\phi)-i\omega\partial_{\phi}]\hat{\Phi},~{}\hat{\Phi}(\mathbf{x},\phi,t)=\left[\begin{array}[]{c}\hat{\varphi}(\mathbf{x},\phi,t)\\ \hat{\varphi}^{\dagger}(\mathbf{x},\phi,t)\end{array}\right].$ (141)
@@
$(z_{n}||z_{m})\equiv\braket{z_{n}}{\sigma_{z}z_{m}}=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi~{}(z_{n}(\phi)|z_{m}(\phi)).$ (142)
The Goldstone-Gibbs modes $z_{a}(\mathbf{x},\phi)$ are given here in terms of $\partial_{a}\Psi_{0}(\mathbf{x},\phi)$, while the Bogoliubov sector is obtained by the adaptation of the Floquet modes of Eq. (121), $z_{\varepsilon,\nu}(\mathbf{x},\phi)\equiv z_{\varepsilon,\nu}(\mathbf{x},\omega t)$. In order to have a complete orthonormal basis in the extended Hilbert space, we promote each mode $z_{n}(\mathbf{x},\phi)$ to a whole set
$z_{n,m}(\mathbf{x},\phi)\equiv z_{n}(\mathbf{x},\phi)e^{im\phi},~{}m\in\mathbb{Z}.$ (143)
Following the prescriptions of Sec. III.2, it is immediate to show that these modes satisfy
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\alpha,m}=m\omega z_{\alpha,m},$
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{A,m}=m\omega z_{A,m}+i\partial_{A}\lambda^{\beta}z_{\beta,m},$ (144)
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\varepsilon,\nu,m}=[\varepsilon+m\omega]z_{\varepsilon,\nu,m},$
@@
$(z_{\varepsilon,\nu,m}||z_{\varepsilon^{\prime},\nu^{\prime},m^{\prime}})$ $=$ $\delta_{\varepsilon\varepsilon^{\prime}}\delta_{\nu\nu^{\prime}}\delta_{mm^{\prime}},$
$(z_{a,m}||z_{\varepsilon,\nu,m^{\prime}})$ $=$ $0,$ (145)
@@
Their quantum amplitude is obtained by $\hat{\gamma}_{n,m}(t)=(z_{n,m}||\hat{\Phi}(t))$. From the Fourier relations
$\hat{\gamma}_{n,m}(t)$ $=$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi~{}e^{-im\phi}\hat{\gamma}_{n}(\phi,t),$ (146)
$\hat{\gamma}_{n}(\phi,t)$$\equiv$ $(z_{n}(\phi)|\hat{\Phi}(\phi,t))=\sum_{m=-\infty}^{\infty}\hat{\gamma}_{n,m}(t)e^{im\phi},$
@@
$\hat{\Phi}(\mathbf{x},\phi,t)$ $=$ $\sum_{m=-\infty}^{\infty}\hat{X}_{m}^{\alpha}(t)z_{\alpha,m}(\mathbf{x},\phi)+\hat{P}_{m}^{A}(t)z_{A,m}(\mathbf{x},\phi)+\sum_{m=-\infty}^{\infty}\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu,m}(t)z_{\varepsilon,\nu,m}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{\varepsilon,\nu.m}(t)\bar{z}_{\varepsilon,\nu,m}(\mathbf{x},\phi)$ (147)
$=$ $\hat{X}^{\alpha}(\phi,t)z_{\alpha}(\mathbf{x},\phi)+\hat{P}^{A}(\phi,t)z_{A}(\mathbf{x},\phi)+\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}(\phi,t)z_{\varepsilon,\nu}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}(\phi,t)\bar{z}_{\varepsilon,\nu}(\mathbf{x},\phi).$
@@
The dynamics for $\hat{\Phi}(\mathbf{x},\phi,t)$ is readily obtained by the relation $i\partial_{t}\hat{\gamma}_{n,m}(t)=([M_{0}(\phi)-i\omega\partial_{\phi}]z_{n,m}||\hat{\Phi}(t))$, finding:
@@
$\hat{\gamma}_{\varepsilon,\nu,m}(t)=\hat{\gamma}_{\varepsilon,\nu,m}e^{-i[\varepsilon+m\omega]t},~{}~{}~{}\hat{P}_{m}^{A}(t)=\hat{P}_{m}^{A}e^{-im\omega t},~{}~{}~{}\hat{X}_{m}^{\alpha}(t)=[\hat{X}_{m}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}_{m}^{B}t]e^{-im\omega t}.$ (148)
When setting $\phi=\phi_{0}+\omega t$, after noticing that the standard time-dependent modes and amplitudes are $z_{n}(\mathbf{x},t)=z_{n}(\mathbf{x},\phi_{0}+\omega t)$ and $\hat{\gamma}_{n}(t)=\hat{\gamma}_{n}(\phi_{0}+\omega t,t)$, we recover the expected time evolution for the field fluctuations:
@@
$\hat{\Phi}(\mathbf{x},t)=\hat{\Phi}(\mathbf{x},\phi_{0}+\omega t,t)=[\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t]z_{\alpha}(\mathbf{x},t)+\hat{P}^{A}z_{A}(\mathbf{x},t)+\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}z_{\varepsilon,\nu}(\mathbf{x},t)e^{-i\varepsilon t}+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}\bar{z}_{\varepsilon,\nu}(\mathbf{x},t)e^{i\varepsilon t}.$ (149)
The crucial difference with respect to Eq. (VI) is that the Gibbs modes there encode a non-trivial trivial dependence in $z_{\phi}$ since
$\partial_{A}\Psi_{0}(\mathbf{x},t)=\partial_{A}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t)+\partial_{\phi}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t)\partial_{A}\omega t.$ (150)
@@
When this dependence is taken into account, as well as the thermodynamic relation between $K$ and $F$, after some tedious algebra one arrives at the correct expansion (149) from the original Floquet modes of Eqs. (120), (121). We note that the above expansion is also valid for driven Floquet systems by excluding the Goldstone-Gibbs pair $\{z_{\phi},z_{F}\}$ and using $\lambda_{A}=\partial_{A}I$ instead of $\lambda_{A}=\partial_{A}E$.
VI.4 Time operator
@@
The linear dependence arises due to the quantum fluctuations of the frequency since this depends on the conserved charges, in analogy to the stationary case [see Eq. (77) and ensuing discussion]. The actual temporal FNG mode, corresponding to proper time translations, is simply obtained as $z_{t}=\omega z_{\phi}$, with an amplitude $\hat{X}^{t}=\hat{X}^{\phi}/\omega$.
On the other hand, the quantum fluctuations $\delta\hat{F}$ of the Floquet charge are those of the more physical generalized Gibbs energy,
@@
In general, no time operator $\hat{T}$ is allowed in Quantum Mechanics since the generator of time translations is the Hamiltonian itself, $[\hat{T},\hat{H}]=-i$, so
$e^{i\hat{T}E_{0}}\hat{H}e^{-i\hat{T}E_{0}}=\hat{H}+E_{0}.$ (154)
for arbitrary value of $E_{0}$. This implies that, if $E$ is an eigenenergy of the Hamiltonian, then $E+E_{0}$ is also an eigenergy. However, Hamiltonians are bounded from below by the ground state energy, leading to a contradiction.
This relation is very similar to that between the particle number and phase operators, which we review following the enlightening discussion of Ref. Mora and Castin (2003). If a phase operator $\hat{\theta}$ exists, then it satisfies $[\hat{N},\hat{\theta}]=i$. But
$e^{i\hat{\theta}N_{0}}\hat{N}e^{-i\hat{\theta}N_{0}}=\hat{N}-N_{0}$ (155)
implies that, if $N$ is an eigenvalue of $\hat{N}$, then $N-N_{0}$ is also. This violates two fundamental properties of the spectrum of the particle number operator: its positive definiteness and its discreteness. Nevertheless, one can define phase fluctuations in situations involving large particle numbers, where one can neglect the discreteness of the spectrum and states with low occupation number, as in the case of a condensate. Analogously, by identifying energy with particle number and time with phase, one can define time fluctuations for large energies, well above the ground state. Actually, within the $(t,\phi)$ formalism, the time operator is a sort of phase operator. This analogy suggests that spontaneous Floquet states should be highly excited states, shifted by a macroscopic energy from the ground state.
VII Time supersolids: CES state
@@
Figure 2: (a) Schematic profile of the sound (solid blue) and flow (dashed red) velocities of the initial homogeneous flowing condensate, $\Psi(x,0)=e^{ivx}$. The arrow represents the attractive delta potential $V(x)=-Z\delta(x)$, suddenly introduced at $t=0$. (b) Dynamical phase diagram for the final state of the system as a function of $(v,Z)$. (c) 2D plot of $|\Psi(x,t)|^{2}$ resulting from the evolution of (a) for $v=0.8$ and $Z=1$. (d) Critical behavior of the CES frequency $\omega$ close to the phase transition along the green lines in (b). The red line represents a fit to a power law. Main panel: Velocity dependence. Inset: Delta-strength dependence.
@@
Figure 3: Snapshots of the normalized density-density correlation function $G(x,x^{\prime},t)$, computed within the Truncated Wigner approximation for a CES state with the same mean-field parameters of Fig. 2c, whose period is $T\approx 12.0$.
We study in this section a specific realization of a spontaneous Floquet state, the so-called CES state de Nova and Sols (2022), which is a solution of a 1D time-dependent GPG equation (117) of the form:
@@
The presence of an inhomogeneous external potential $V(x)\neq 0$ explicitly breaks invariance under spatial translations. This guarantees the genuine character of the spontaneous symmetry-breaking of time translation symmetry, avoiding any misidentification of time crystalline behavior as that discussed after Eq. (107). Although the CES state was originally identified in an analogue gravity context de Nova et al. (2016, 2021), similar scenarios involving soliton emission and periodic self-oscillations are well-known in the literature Hakim (1997); Pavloff (2002); Engels and Atherton (2007); Nguyen et al. (2017); Sels and Demler (2020); Tamura et al. (2023).
In analogy to the case of Sec. IV, the CES state spontaneously breaks $U(1)$ and time-translation symmetries, combining superfluidity with time crystalline behavior. This represents the temporal analogue of a supersolid, which has been denoted as a time supersolid and was recently observed in a magnonic BEC Autti et al. (2018). We remark that, while strictly speaking the cnoidal wave is not a supersolid because it is not at equilibrium Martone et al. (2021), time crystals are necessarily out-of-equilibrium states as a consequence of the celebrated no-go theorem Watanabe and Oshikawa (2015). In the presence of external periodic driving, self-interacting condensates governed by the GP equation can also give rise to discrete time crystals Sacha (2015).
Due to the phase and time-translational invariance, if $\Psi_{0}(x,t)$ is a solution of Eq. (156), then $e^{-i\theta}\Psi_{0}(x,t+t_{0})$ is also a solution. In the $(t,\phi)$ formalism, where $\Psi_{0}(x,t+t_{0})=\Psi_{0}(x,\omega t+\phi_{0})$, the CES state is characterized by the vectors $\alpha=(\theta,\phi)$, $T_{\alpha}=(1,i\partial_{\phi})$, $Q_{\alpha}=(N,F)$, and $\lambda^{\alpha}=(\mu,\omega)$. The Floquet-Goldstone-Gibbs modes are then
$z_{\theta}(x,\phi)$ $=$ $\left[\begin{array}[]{r}-i\Psi_{0}(x,\phi)\\ i\Psi_{0}^{*}(x,\phi)\end{array}\right],~{}z_{\phi}(x,\phi)=\left[\begin{array}[]{l}\partial_{\phi}\Psi_{0}(x,\phi)\\ \partial_{\phi}\Psi_{0}^{*}(x,\phi)\end{array}\right],$ (161)
@@
The CES state satisfies typical time-crystal criteria of robustness, independence from the initial condition, and universality de Nova and Sols (2022). In this work, we focus on a specific realization in which a localized attractive delta barrier $V(x)=-Z\delta(x)$ is quenched at $t=0$ within a subsonic homogeneous condensate flowing with velocity $v$, described by a GP wavefunction $\Psi(x,0)=e^{ivx}$ (in our choice of units, the initial condensate density is then $n_{0}$). A schematic depiction of the initial setup is displayed in Fig. 2a. The quench in the external potential induces a deterministic dynamics in the condensate, numerically computed by integrating the time-dependent GP equation (22). The asymptotic behavior of the system is described by a dynamical phase diagram which is solely function of $v,Z$, Fig. 2b, exhibiting only two possible final states: the nonlinear ground state (GS) Michel and Parentani (2013) or the CES state. Specifically, for initial velocities larger than the critical velocity $v>v_{c}(Z)$, or, equivalently, for barrier amplitudes larger than the critical amplitude $Z>Z_{c}(v)$, the GP condensate wavefunction asymptotically approaches the CES state along the lines of Eq. (118),
@@
where it oscillates periodically. This periodic behavior is globally displayed both upstream and downstream, as seen in Fig. 2c, and not just in the traveling cnoidal wave downstream (whose time periodicity is rather trivial, as previously argued). We note that the CES state is a macroscopically excited state, something expected from the discussion at the end of Sec. VI.4, since the initial state is shifted by an energy $\Delta E=Nv^{2}/2$ with respect to a condensate at rest.
The GS/CES phase diagram is an example of dynamical phase transition Moeckel and Kehrein (2008); Sciolla and Biroli (2010); Lang et al. (2018), where the GS is the symmetry-unbroken phase, with continuous time translation symmetry, while the CES state is the time-crystalline phase, with discrete time translation symmetry. The oscillation frequency $\omega$ exhibits a critical behavior close to the phase boundary, where the critical exponents for $v,Z$ (obtained from a fit in Fig. 2d) are both approximately $\gamma\simeq\delta\simeq 0.50$, in agreement with those for a square well de Nova and Sols (2022), potentially suggesting a possible analytical derivation.

### v4

@@ 6.3 Quantization
which is the quantum version of the time-dependent Floquet-GPG equation. A similar equation can be derived for a conventional Floquet system, where the frequency is fixed by the external driving provided by the periodic Hamiltonian $\hat{H}(\phi)$. Once the dynamics in the $(t,\phi)$ formalism is computed, the actual field operator is retrieved as $\hat{\Psi}(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\hat{\Psi}(\mathbf{x},\phi_{0}+\omega t,t)$.
After expanding the field operator around the stationary Floquet state, $\hat{\Psi}(\mathbf{x},\phi,t)=\Psi_{0}(\mathbf{x},\phi)+\hat{\varphi}(\mathbf{x},\phi,t)$, we obtain at linear order from Eq. (145) that
$i\partial_{t}\hat{\Phi}$ $=$ $[M_{0}(\phi)-i\omega\partial_{\phi}]\hat{\Phi}$,
$\hat{\Phi}(\mathbf{x},\phi,t)$ $=$ $\left[\begin{array}[]{c}\hat{\varphi}(\mathbf{x},\phi,t)\\ \hat{\varphi}^{\dagger}(\mathbf{x},\phi,t)\end{array}\right].$ (148)
@@
$(z||z^{\prime})\equiv\braket{z\|\sigma_{z}z^{\prime}}=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\leavevmode\nobreak\ (z(\phi)|z^{\prime}(\phi)).$ (149)
The Goldstone-Gibbs modes $z_{a}(\mathbf{x},\phi)$ are given here in terms of $\partial_{a}\Psi_{0}(\mathbf{x},\phi)$, while the Bogoliubov sector $z_{n}(\mathbf{x},\phi)$ is obtained by the adaptation of the Floquet modes of Eq. (126), $z_{n}(\mathbf{x},t)\equiv z_{n}(\mathbf{x},\phi_{0}+\omega t)$. In order to have a complete orthonormal basis in the extended Hilbert space, we promote each spinor $z(\mathbf{x},\phi)$ to a whole discrete set
$z_{m}(\mathbf{x},\phi)\equiv z(\mathbf{x},\phi)e^{im\phi},\leavevmode\nobreak\ m\in\mathbb{Z}.$ (150)
Following the prescriptions of Secs. 3.2, 3.3, it is immediate to show that these modes satisfy
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\alpha,m}=m\omega z_{\alpha,m},$ (151)
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{A,m}=m\omega z_{A,m}+i\partial_{A}\lambda^{\beta}z_{\beta,m},$
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{n,m}=[\varepsilon_{n}+m\omega]z_{n,m},$
@@
$(z_{n,m}||z_{n^{\prime},m^{\prime}})$ $=$ $\delta_{nn^{\prime}}\delta_{mm^{\prime}},$ (152)
$(z_{a,m}||z_{n,m^{\prime}})$ $=$ $0,$
@@
Notice that here the Berry-Gibbs curvature (56) is evaluated using the extended scalar product (135).
Quantum amplitudes are now obtained by $\hat{\gamma}_{m}(t)=(z_{m}||\hat{\Phi}(t))$. From the Fourier relations
$\hat{\gamma}_{m}(t)$ $=$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\leavevmode\nobreak\ e^{-im\phi}\hat{\gamma}(\phi,t),$ (153)
$\hat{\gamma}(\phi,t)$ $\equiv$ $(z(\phi)|\hat{\Phi}(\phi,t))=\sum_{m=-\infty}^{\infty}\hat{\gamma}_{m}(t)e^{im\phi},$
@@
$\hat{\Phi}(\mathbf{x},\phi,t)$ $=\sum_{m=-\infty}^{\infty}\hat{X}_{m}^{\alpha}(t)z_{\alpha,m}(\mathbf{x},\phi)+\hat{P}_{m}^{A}(t)z_{A,m}(\mathbf{x},\phi)+\sum_{m=-\infty}^{\infty}\sum_{n}\hat{\gamma}_{n,m}(t)z_{n,m}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{n.m}(t)\bar{z}_{n,m}(\mathbf{x},\phi)$
$=\hat{X}^{\alpha}(\phi,t)z_{\alpha}(\mathbf{x},\phi)+\hat{P}^{A}(\phi,t)z_{A}(\mathbf{x},\phi)+\sum_{n}\hat{\gamma}_{n}(\phi,t)z_{n}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{n}(\phi,t)\bar{z}_{n}(\mathbf{x},\phi).$ (154)
@@
The dynamics for $\hat{\Phi}(\mathbf{x},\phi,t)$ is readily obtained by the relation $i\partial_{t}\hat{\gamma}_{m}(t)=([M_{0}(\phi)-i\omega\partial_{\phi}]z_{m}||\hat{\Phi}(t))$, finding:
@@
$\hat{\gamma}_{n,m}(t)=\hat{\gamma}_{n,m}e^{-i[\varepsilon_{n}+m\omega]t},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \hat{P}_{m}^{A}(t)=\hat{P}_{m}^{A}e^{-im\omega t},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \hat{X}_{m}^{\alpha}(t)=[\hat{X}_{m}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}_{m}^{B}t]e^{-im\omega t}.$ (155)
Inverting the Fourier transformation yields:
@@
$\hat{\gamma}_{n}(\phi,t)=\hat{\gamma}_{n}(\phi-\omega t,0)e^{-i\varepsilon_{n}t},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \hat{P}^{A}(\phi,t)=\hat{P}^{A}(\phi-\omega t,0),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \hat{X}^{\alpha}(\phi,t)=\hat{X}^{\alpha}(\phi-\omega t,0)+\partial_{A}\lambda_{B}\hat{P}^{B}(\phi-\omega t,0)t.$ (156)
When setting $\phi=\phi_{0}+\omega t$, after noticing that the standard time-dependent spinors and amplitudes are $z(\mathbf{x},t)=z(\mathbf{x},\phi_{0}+\omega t)$ and $\hat{\gamma}(t)=\hat{\gamma}(\phi_{0}+\omega t,t)$, with $\hat{\gamma}\equiv\hat{\gamma}(t=0)=\hat{\gamma}(\phi_{0},0)$ their initial values, we recover the expected time evolution for the field fluctuations:
$\hat{\Phi}(\mathbf{x},t)=\hat{\Phi}(\mathbf{x},\phi_{0}+\omega t,t)=[\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t]z_{\alpha}(\mathbf{x},t)+\hat{P}^{A}z_{A}(\mathbf{x},t)+\sum_{n}\hat{\gamma}_{n}z_{n}(\mathbf{x},t)e^{-i\varepsilon_{n}t}+\hat{\gamma}^{\dagger}_{n}\bar{z}_{n}(\mathbf{x},t)e^{i\varepsilon_{n}t}.$ (157)
Formally, this is the same expansion of Eqs. (6), (6), which in turn presented the same structure of the stationary case, Eq. (3.4). The crucial difference is that now we use $(\phi,\omega,F)$ instead of $(t,1,K)$, so the Lagrange multiplier $\lambda^{\phi}=\omega=\partial_{F}E$ does lead to the correct (thermo)dynamics. The failure of the original time-dependent approach is that the Gibbs modes there encode a non-trivial dependence in $z_{\phi}$ since
$\partial_{A}\Psi_{0}(\mathbf{x},t)$ $=$ $\partial_{A}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t)$
$+$ $\partial_{\phi}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t)\partial_{A}\omega t.$
@@
When this dependence is taken into account, as well as the thermodynamic relation between $K$ and $F$, after some tedious algebra one arrives at the correct expansion (157) from the original Floquet modes of Eqs. (125), (126).
We note that the above expansion is also valid for driven Floquet systems by excluding the Goldstone-Gibbs pair $\{z_{\phi},z_{F}\}$ and using $\lambda_{A}=\partial_{A}I$ instead of $\lambda_{A}=\partial_{A}E$.
In summary, the $(t,\phi)$ formalism allows to treat Floquet states as stationary states by replacing time-independent wavefunctions by periodic wavefunctions, the generalized Gibbs energy $K$ by the generalized Floquet-Gibbs energy $\Lambda=K-\omega F$, energies $\epsilon$ by quasi-energies $\varepsilon$, and conserved charges by their angular averages:
$\Psi_{0}(\mathbf{x}),z(\mathbf{x})$ $\longrightarrow$ $\Psi_{0}(\mathbf{x},\phi),z(\mathbf{x},\phi),$ (159)
$K$ $\longrightarrow$ $\Lambda,$
$\epsilon$ $\longrightarrow$ $\varepsilon,$
$Q_{\alpha}$ $\longrightarrow$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\leavevmode\nobreak\ Q_{\alpha}(\phi).$
For spontaneous Floquet states, one also needs to add the Floquet charge $F$, with no stationary counterpart, while for driven Floquet states the energy $E$ is replaced by the Floquet enthalpy $I$ as the correct thermodynamic potential.
6.4 Time operator
@@
The linear time dependence arises due to the quantum fluctuations of the frequency since this depends on the conserved charges, in analogy to the stationary case [see Eq. (81) and ensuing discussion]. The actual temporal FNG mode, corresponding to proper time translations, is simply obtained as $z_{t}=\omega z_{\phi}$, with an amplitude $\hat{X}^{t}=\hat{X}^{\phi}/\omega$.
On the other hand, the quantum fluctuations $\delta\hat{F}$ of the Floquet charge are proportional to those of the more physical generalized Gibbs energy,
@@
In general, no time operator $\hat{T}$ is allowed in Quantum Mechanics since the generator of time translations is the Hamiltonian itself, $\hat{T},\hat{H}]=-i$, so
$e^{-i\hat{T}E_{0}}\hat{H}e^{i\hat{T}E_{0}}=\hat{H}-E_{0}$ (163)
for arbitrary value of $E_{0}$. This implies that, if $E$ is an eigenenergy of the Hamiltonian, then $E-E_{0}$ is also an eigenergy. However, Hamiltonians are bounded from below by the ground-state energy, leading to a contradiction.
This relation is very similar to that between the particle number and phase operators, which we review following the enlightening discussion of Ref. [87]. If a phase operator $\hat{\theta}$ exists, then it satisfies $\hat{N},\hat{\theta}]=i$. But
$e^{-i\hat{\theta}N_{0}}\hat{N}e^{i\hat{\theta}N_{0}}=\hat{N}-N_{0}$ (164)
implies that, if $N$ is an eigenvalue of $\hat{N}$, then $N-N_{0}$ is also. This violates two fundamental properties of the spectrum of the particle-number operator: its positive definiteness and its discreteness. Nevertheless, one can define phase fluctuations in situations involving large particle numbers, where one can neglect the discreteness of the spectrum and states with low occupation number, as in the case of a condensate. Analogously, by identifying energy with particle number and time with phase, one can define time fluctuations for large energies, well above the ground state. Actually, within the $(t,\phi)$ formalism, the time operator is a sort of phase operator. This analogy suggests that spontaneous Floquet states should be highly excited states, shifted by a macroscopic energy from the ground state.
7 Time supersolids: CES state
@@
Figure 3: (a) Schematic profile of the sound (solid blue) and flow (dashed red) velocities of the initial homogeneous flowing condensate, $\Psi(x,0)=e^{ivx}$. The arrow represents the attractive delta potential $V(x)=-Z\delta(x)$, suddenly introduced at $t=0$. (b) Dynamical phase diagram for the final state of the system as a function of $(Z,v)$. (c) 2D plot of $|\Psi(x,t)|^{2}$ resulting from the evolution of (a) for $v=0.8$ and $Z=1$. (d) Critical behavior of the CES frequency $\omega$ close to the phase transition along the green lines in (b). The red line represents a fit to a power law. Main panel: Velocity dependence. Inset: Delta-strength dependence.
@@
Figure 4: Snapshots of the normalized density-density correlation function $G(x,x^{\prime},t)$, computed using the Truncated Wigner method for a CES state with the same mean-field parameters of Fig. 3c, whose period is $T\approx 12.0$.
We study in this section a specific realization of spontaneous Floquet state, the so-called CES state [43], which is a periodic solution of a 1D time-dependent GPG equation (122) of the form:
@@
The presence of an inhomogeneous external potential $V(x)\neq 0$ explicitly breaks invariance under spatial translations. This guarantees the genuine character of the spontaneous symmetry-breaking of time-translation symmetry, avoiding any misidentification of time crystalline behavior as that discussed after Eq. (112). Although the CES state was originally identified in an analogue gravity context [44, 45], similar scenarios involving soliton emission and periodic self-oscillations are well-known in the literature [88, 89, 90, 91, 92, 93].
In analogy to the case of Sec. 4, the CES state spontaneously breaks $U(1)$ and time-translation symmetries, combining superfluidity with time crystalline behavior. This represents the temporal analogue of a supersolid, which has been denoted as time supersolid and was recently observed in a magnonic BEC [27]. We remark that, while strictly speaking the cnoidal wave is not a supersolid because it is not at equilibrium [57], time crystals are necessarily out-of-equilibrium states as a consequence of their celebrated no-go theorem [18]. In the presence of external periodic driving, self-interacting condensates governed by the GP equation can also give rise to discrete time crystals [32].
Due to the phase and time-translation invariance, if $\Psi_{0}(x,t)$ is a solution of Eq. (165), then $e^{-i\theta}\Psi_{0}(x,t+t_{0})$ is also a solution. In the $(t,\phi)$ formalism, where $\Psi_{0}(x,t+t_{0})=\Psi_{0}(x,\omega t+\phi_{0})$, the simultaneous symmetry breaking in the CES state is characterized by the vectors $\boldsymbol{\alpha}=(\theta,\phi)$, $\mathbf{T}=(1,i\partial_{\phi})$, $\mathbf{Q}=(N,F)$, and $\boldsymbol{\lambda}=(\mu,\omega)$. The Goldstone-Gibbs modes are then
$z_{\theta}(x,\phi)$ $=\left[\begin{array}[]{r}-i\Psi_{0}(x,\phi)\\ i\Psi_{0}^{*}(x,\phi)\end{array}\right],\leavevmode\nobreak\ z_{\phi}(x,\phi)=\left[\begin{array}[]{l}\partial_{\phi}\Psi_{0}(x,\phi)\\ \partial_{\phi}\Psi_{0}^{*}(x,\phi)\end{array}\right],$ (170)
@@
The CES state satisfies typical time-crystal criteria of robustness, independence from the initial condition, and universality [43]. In this work, we focus on a specific realization in which a localized attractive delta barrier $V(x)=-Z\delta(x)$ is quenched at $t=0$ within a subsonic homogeneous condensate flowing with velocity $v$, described by a GP wavefunction $\Psi(x,0)=e^{ivx}$ (in our choice of units, the initial condensate density is then $n_{0}$). A schematic depiction of the initial setup is displayed in Fig. 3a. The quench in the external potential induces a deterministic dynamics in the condensate, numerically computed by integrating the time-dependent GP equation (21). The asymptotic behavior of the system is described by a dynamical phase diagram which is solely function of $(Z,v)$, Fig. 3b, exhibiting only two possible final states: the nonlinear ground state (GS) [94] or the CES state. Specifically, for initial velocities larger than the critical velocity $v>v_{c}(Z)$, or, equivalently, for barrier amplitudes larger than the critical amplitude $Z>Z_{c}(v)$, the GP wavefunction asymptotically approaches the CES state along the lines of Eq. (123),
@@
where it oscillates periodically. This periodic behavior is globally displayed both upstream and downstream, as seen in Fig. 3c, and not just in the traveling cnoidal wave downstream (whose time periodicity is rather trivial, as previously argued). We note that the CES state is a macroscopically excited state, something expected from the discussion at the end of Sec. 6.4, since the initial state is shifted by an energy $\Delta E=Nv^{2}/2$ with respect to a condensate at rest.
The GS/CES phase diagram is an example of dynamical phase transition [95, 96, 97], where the GS is the symmetry-unbroken phase, with continuous time-translation symmetry, while the CES state is the time-crystalline phase, with discrete time-translation symmetry. The oscillation frequency $\omega$ exhibits a critical behavior close to the phase boundary, where the critical exponents $\delta_{v},\delta_{Z}$ for $v,Z$ (obtained from a fit in Fig. 3d) are both approximately $\delta_{v}\simeq\delta_{Z}\simeq 0.50$, in agreement with those for a square well [43], potentially suggesting a possible analytical derivation. The main advantage of using a delta barrier with respect to a finite square well is that the transient towards the CES state is shorter, reducing the numerical cost of the simulations.